What downward force is exerted on a small piston with a radius of 2.0 cm would lift a 2000.0 kg car if the large piston has a radius of 12.0 cm?


1
Expert's answer
2021-10-24T18:25:17-0400

"\text{for small piston:}"

"r_1 = 2cm"

"S_1=\pi r_1^2= 4\pi \ cm^2"

"p_1 =\frac{F_1}{S_1}"

"\text{\text{for big piston:}}"

"g = 9.81\frac{m}{s^2}"

"m = 2000kg"

"F_2= mg = 2000*9.81=19620N"

"r_2=12cm"

"S_2= \pi r_2^2= 144\pi\ cm^2"

"p_2 = \frac{F_2}{S_2}"

"p_1= p_2"

"F_1= S_1\frac{F_2}{S_2}=4\pi*\frac{19620}{144\pi}=545N"


"\text{Answer: }545N"